OpenStudy (anonymous):

really hard question: draw the graph(sketch) of (y^2-1)sin(cosh(x)^3y)=x^2-4

OpenStudy (anonymous):

if you're just interested of what it looks like: http://www.wolframalpha.com/input/?i=plot+%28y%5E2-1%29sin%28cosh%28x%29%5E3y%29%3Dx%5E2-4

OpenStudy (anonymous):

looks nice, how to mentally draw it?

OpenStudy (anonymous):

Sorry, my input was wrong. I think this is what you mean: http://www.wolframalpha.com/input/?i=plot+%28y%5E2-1%29sin%28cosh%28x%29%5E%283y%29%29%3Dx%5E2-4

OpenStudy (anonymous):

\[(y^2-1)\sin(\cosh(x)^3y)=x^2-4\] \[y^2=x^2-4+\sin(\cosh(x)^3y)\] y^2=x^2-4 is the equation for a hyperbola ( http://www.cs.cornell.edu/w8/~andru/cgi-bin/relplot/plot.pl), so this is just a pimped up hyperbola that varies crazily for x and y.

OpenStudy (anonymous):

Well, that sin(cosh) mess is going to vary between 1 and -1, maybe that helps a bit....